Express Distribution markets CDs of the performing artist Fishe. At the beginning of October, Express had in beginning inventory 5,840 of Fishe’s CDs with a unit cost of $6. During October, Express made the following purchases of Fishe’s CDs.

Oct. 3 7,300 @ $7 Oct. 19 8,760 @ $10
Oct. 9 10,220 @ $8 Oct. 25 11,680 @ $11


During October, 31,828 units were sold. Express uses a periodic inventory system.

Determine (1) the ending inventory and (2) the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). (Round answers to 0 decimal places, e.g. 1,255.)

FIFO

LIFO

AVERAGE-COST

The ending inventory $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

The cost of goods sold $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer
Answer #1
Units Unit cost Total
Beg. Inventory 5840 6 35040
Oct. 3 7300 7 51100
Oct. 9 10220 8 81760
Oct. 19 8760 10 87600
Oct. 25 11680 11 128480
Total 43800 383980
Average cost per unit 8.7667 =383980/43800
Ending inventory units 11972 =43800-31828
The ending inventory:
FIFO 131400 =(11680*11)+(11972-11680)*10
LIFO 77964 =(5840*6)+(11972-5840)*7
Average cost 104955 =11972*8.7667
The cost of goods sold:
FIFO 252580 =383980-131400
LIFO 306016 =383980-77964
Average cost 279025 =383980-104955